An 8.0 gram sample of CO2 contains moles of oxygen atoms A. 0.18 B. 0.36 C. 1.09 x 1023 D. 2.19 x 1023
Answer #1

ANS. B) 0.36

molar mass of CO2= 12 + 16 x 2 = 44g
44g of CO2 gives 1 mol of CO2
8g of CO2 gives 8/44 mol of CO2
1 mol of CO2 has 2 mol of O2  atoms
8/44 mol of CO2 has 16/44 mol of O2 atoms

so, 16/44 mol of O2 atoms = 0.36 moles of O2 atoms

/* PLEASE UPVOTE */